The manufacturer of Brand A automobile tires claims that its tire can save 110 gallons of fuel over 55,000 miles of driving, as compared to a popular competitor (Brand B). If gasoline costs $4.00 per gallon, how much per mile driven does this tire save the customer (Brand A versus Brand B)?
Answer #1

Answer : For Brand A;

55,000 miles of driving saves 110 gallons of fuel

So, 1 mile of driving saves = 110 / 55,000 = 0.002 gallon of fuel.

Given, 1 gallon's cost is $4

So, 0.002 gallon's cost = 4 * 0.002 = $0.008

Therefore, Brand A automobile tires saves $0.008 for customer's per mile driven in compared to Brand B automobile tires.